1. What are the vector components of the following: Round off your answer to two decimal numbers.

a. G = 10m, 1500°

b. L = 10m, 900°

c. A = 10m, 3600°

d. D = 10m, 360° N of W


1
Expert's answer
2021-09-26T08:40:54-0400

"\text{Part a.}\\
G_x=G\cos\theta=-8.66\text{ m},\\
G_y=G\sin\theta=5\text{ m}.\\\space\\
\text{Part b.}\\
L_x=L\cos\theta=0\text{ m},\\
L_y=L\sin\theta=10\text{ m}.\\\space\\
\text{Part c.}\\
A_x=A\cos\theta=10\text{ m},\\
A_y=A\sin\theta=0\text{ m}.\\\space\\
\text{Part d.}\\
D_x=D\cos\theta=-10\text{ m},\\
D_y=D\sin\theta=0\text{ m}."